### v1

# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

Abstract
The recent ATOMKI experiments provided evidence pointing towards the existence of an X17 boson in the anomalous nuclear transitions of Beryllium-8, Helium-4, and Carbon-12. In this work, we consider X17 boson contributions to the previously measured $D$ meson decays which include $D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-}$, $D_{s}^{*+}\rightarrow D_{s}^{+}\gamma$, $D^{*0}\rightarrow D^{0}e^{+}e^{-}$, and $D^{*0}\rightarrow D^{0}\gamma$, as well as the measured decays of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$. Using the data of the above meson decays, we perform a fitting to the coupling parameters $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ by treating the couplings $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent from each other rather than assuming the generation universality $\varepsilon_{u}=\varepsilon_{c}$. It is found that the above fitting renders $|\varepsilon_{c}|=3\times 10^{-4}$, $|\varepsilon_{s}|=4\times 10^{-5}$ and a huge magnitude for $\varepsilon_{u}$, which is in serious tension with $\left|\varepsilon_{u}\right|$ determined from ATOMKI measurements.
I Introduction
In the search for new physics beyond the standard model, the study of low energy processes is an effective avenue besides those focusing on either high energy or precision frontiers. There has been a persistent anomaly in the measured angular distribution of $e^{+}e^{-}$ pairs in the 18.15 MeV decay of the excited state 8Be* by the ATOMKI collaboration over the last few years Krasznahorkay:2015iga ; Krasznahorkay:2017gwn ; Krasznahorkay:2017qfd ; Krasznahorkay:2019lgi . The simplest way to account for the data is to introduce a new X17 boson with a mass in the range of approximately 16 to 17 MeV. The de-excitation of ${}^{8}{\rm Be}^{*}$ is assumed to emit a X17 boson which then decays to the $e^{+}e^{-}$ pair as detected by the above experiments. It is desirable to confirm the existence of X17 boson which undoubtedly leads to new physics beyond the Standard Model (SM).
The results reported in Krasznahorkay:2015iga gives $m_{X}=16.70\pm 0.35({\rm stat})\pm 0.5({\rm syst})~{}\rm{MeV/c^{2}}$ and the ratio $\Gamma(^{8}{\rm Be}(18.15)\rightarrow~{}^{8}{\rm Be}+{X})\times{\rm BR}({X}\rightarrow e^{+}e^{-})/\Gamma(^{8}{\rm Be}(18.15)\rightarrow~{}^{8}{\rm Be}+\gamma)\equiv R_{X}=5.8\times 10^{-6}$, with a statistical significance of $6.8\sigma$. Later analysis from more 8Be experiments gave the updated results, $m_{X}=17.01\pm 0.16~{}\rm{MeV/c^{2}}$ and $R_{X}=(6\pm 1)\times 10^{-6}$ Krasznahorkay:2019lgi . This is also the most recent data for the proposed X17 boson resulting from the 8Be anomaly, regardless of the significant anomalies observed in the case of direct proton capture Sas:2022pgm .
In recent years, there were further studies that suggest potential connections between X17 boson and additional anomalies observed in the decays of excited 4He Firak:2020eil ; Krasznahorkay:2019lyl ; Krasznahorkay:2021joi and 12Krasznahorkay:2022pxs . These anomalies, like the 8Be case, exhibit deviations from the expected decay distributions. The masses of the proposed particle, as inferred from the anomalies observed in excited 4He and 12C decays, are reported as $16.94\pm 0.12({\rm stat})\pm 0.21({\rm syst})~{}\rm{MeV/c^{2}}$ Krasznahorkay:2021joi and $17.03\pm 0.11({\rm stat})\pm 0.2({\rm syst})~{}\rm{MeV/c^{2}}$ Krasznahorkay:2022pxs , respectively. It is remarkable that the mass values for the proposed particle as extracted from various experiments involving different nuclear systems are rather similar. This hints at a unique particle that can account for decay anomalies of excited 8Be, 4He, and 12C [^1] We note that there existed other experiments that perform searches for X17 boson as well. The search by VNU, Vietnam Anh:2024req confirmed the ${}^{8}{\rm Be}^{*}$ decay anomaly while the search by MEG II collaboration at PSI did not observe the resonance MEGII:2024urz . .
There existed studies proposing vector couplings between X17 boson and fermions as the favorable scenario for addressing the above decay anomalies Feng:2016jff ; Feng:2016ysn ; Pulice:2019xel ; Feng:2020mbt ; Nomura:2020kcw . The scenario for axial-vector or more generally the mixture of vector and axial-vector couplings were also proposed  Feng:2020mbt ; Kozaczuk:2016nma ; DelleRose:2017xil ; DelleRose:2018eic ; DelleRose:2018pgm ; Seto:2020jal . There were also proposals that the above decay anomalies can be resolved or mitigated by resorting to various unaccounted effects within SM  Zhang:2017zap ; Aleksejevs:2021zjw ; Wong:2022jeq . In this article we adopt the proposal that X17 is a vector boson. We begin by writing the couplings between X17 boson and quarks of various flavors as
${\cal L}_{X(Q,q)}=e\varepsilon_{Q}X_{\mu}\left(\overline{Q}\gamma^{\mu}Q\right)+e\varepsilon_{q}X_{\mu}\left(\overline{q}\gamma^{\mu}q\right),$ (1)
where $q$ stands for $u$, $d$ and $s$ quarks, while $Q$ represents heavy quarks. From the data of ${}^{8}{\rm Be}$ anomaly, $R_{X}=5.8\times 10^{-6}$, with the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, it has been established that Feng:2016jff ; Fornal:2017msy $\varepsilon_{u}\simeq\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}\simeq\mp 7.4\times 10^{-3}$, where the null result of $\pi^{0}\to X\gamma$ search by NA48/2 NA482:2015wmo has been used. Further refined analysis takes into account isospin effects and their mixings Feng:2016ysn . Such an analysis considers three distinct isospin scenarios: no isospin effects, isospin mixing, and isospin mixing & breaking. Assuming that X17 is a vector boson, a statistical $\chi^{2}$ test based upon data from anomalous transitions observed in excited ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$ has been performed in Denton:2023gat . Utilizing eleven measurements that include angular and width data from the anomalous transitions in ATOMKI experiments, and the null result of $\pi^{0}\to X\gamma$ search NA482:2015wmo under the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, the above $\chi^{2}$ test determines the ranges of three parameters, $m_{X},\varepsilon_{n},\,{\rm and}\,\varepsilon_{p}$ where the latter two are related to the quark-level couplings by $\varepsilon_{p}=2\varepsilon_{u}+\varepsilon_{d}$ and $\varepsilon_{n}=2\varepsilon_{d}+\varepsilon_{u}$. It was found that  Denton:2023gat $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ and $\left|\varepsilon_{d}\right|\simeq(2.5-2.9)\times 10^{-3}$ with $\varepsilon_{u}$ and $\varepsilon_{d}$ in opposite signs. Specifically the most favored parameter set is $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ if isospin effects are not considered. With both isospin mixing and breaking effects considered, the most favored parameter set becomes $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. It is noteworthy that these fitted values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ are much smaller than those determined earlier.
Given the hypothesis of X17 boson to account for the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$ excited states, it is of interest to test this hypothesis on heavy hadrons which involve the couplings of X17 boson to the second and the third generation of quarks, in addition to the couplings $\varepsilon_{u}$ and $\varepsilon_{d}$. In fact, ${}^{8}{\rm Be}$ anomaly already motivated the study of X17 boson effect to $K^{+}\to\mu^{+}\nu_{\mu}e^{+}e^{-}$ decay through the interaction/decay chain, $K^{+}\to\mu^{+}\nu_{\mu}X\to\mu^{+}\nu_{\mu}e^{+}e^{-}$ Chiang:2016cyf with the assumption $\varepsilon_{s}=\varepsilon_{d}$. It was also proposed that X17 boson can be searched by BESIII or Belle II by the decay chain $J/\Psi\to\eta_{c}X^{*}\to\eta_{c}e^{+}e^{-}$ Ban:2020uii , which is an excess to the regular decay chain $J/\Psi\to\eta_{c}\gamma^{*}\to\eta_{c}e^{+}e^{-}$. Furthermore, effects of X17 boson to decays of $D^{*(+,0)}$, $D^{*+}_{s}$, and the corresponding $B$ mesons were also studied in Castro:2021gdf . In the latter work, it was stated that the theoretical prediction of $R_{D_{s}^{*+}}\equiv\Gamma(D_{s}^{*+}\rightarrow D^{+}_{s}e^{+}e^{-})/\Gamma(D_{s}^{*+}\rightarrow D^{+}_{s}\gamma)$ with the contribution $D_{s}^{*+}\to D^{+}_{s}+X\to D^{+}_{s}+e^{+}+e^{-}$ included in the numerator is consistent with the CLEO measurement $R_{D_{s}^{*+}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . We note that the authors of Ref. Castro:2021gdf calculated the X17 contribution to $R_{D_{s}^{*+}}$ by taking $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$  Feng:2016jff and assuming $\varepsilon_{s}=\varepsilon_{d}$, and $\varepsilon_{c}=\varepsilon_{u}$. Furthermore the signals due to X17 boson in other decay modes of $D^{*}$ and $B^{*}$ mesons were predicted.
In this article, we shall revisit the contributions of X17 boson to $D^{*0}$ and $D^{*+}_{s}$ meson decays since there existed a measurement on $R_{D^{*0}}\equiv\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ after the publication of Castro:2021gdf . The measurement gives $R_{D^{*0}}=(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq . Furthermore, we are motivated to tackle this issue because the updated values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ Denton:2023gat are much smaller than those determined earlier Feng:2016jff . We found that the calculation on $R_{D_{s}^{*+}}$ by Castro:2021gdf is incorrect. After correcting the calculation, we found that the prediction of $R_{D_{s}^{*+}}$ is consistent with CLEO data only for the updated values of $\varepsilon_{c}\equiv\varepsilon_{u}$ and $\varepsilon_{s}\equiv\varepsilon_{d}$. On the other hand, the data on $R_{D^{*0}}$ requires a much larger $\varepsilon_{u}$ with the assumption $\varepsilon_{c}=\varepsilon_{u}$. It is obvious that the updated ranges of $\varepsilon_{u}$ and $\varepsilon_{d}$ with the assumptions $\varepsilon_{c}=\varepsilon_{u}$, and $\varepsilon_{s}=\varepsilon_{d}$ cannot simultaneously account for $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$. To ease the tension, we note that the contribution of X17 boson to $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ with $\varepsilon_{c}$ and $\varepsilon_{s}$ in opposite signs as we shall see in Fig. 3 later. Hence one can account for both $R_{D^{*0}}$ and $R_{D_{s}^{*+}}$ by raising both $\left|\varepsilon_{c}\right|$ and $\left|\varepsilon_{s}\right|$ while maintaining $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ fixed. On the other hand, this implies that the assumption $\varepsilon_{s}=\varepsilon_{d}$ is no longer appropriate. In view of this, one may lift the restriction $\varepsilon_{c}=\varepsilon_{u}$ as well. Finally, we note that the enhancement of both $\varepsilon_{c}$ and $\varepsilon_{s}$ should impact the decays of charmonium and $\phi$ meson. Therefore we also include $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$ in our analysis. We shall perform fittings to the coupling strengths $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ with the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}\equiv\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ BESIII:2022jde , and $R_{\phi}\equiv\Gamma(\phi\rightarrow\eta e^{+}e^{-})/\Gamma(\phi\rightarrow\eta\gamma)$ Achasov:2000ne ; CMD-2:2000qla ; KLOE-2:2014hfk . It is important to compare the favored value of $\varepsilon_{u}$ in this fitting to that obtained from the decay anomalies of excited 8Be, 4He, and 12C. Here we fix the mass of X17 boson at $16.85$ MeV which is the best-fit value of $m_{X}$ obtained in Denton:2023gat . We adopt this simplification since various decay rates considered here are rather insensitive to $m_{X}$ over its allowed range.
This article is organized as follows. In Sec. II, we review previous studies on the X17 contributions to heavy meson decays $H^{*}\to He^{+}e^{-}$ normalized by $H^{*}\to H\gamma$ with $H^{*}$ representing $D^{*0}$, $D^{*+}$, and $D^{*+}_{s}$ vector mesons and $H$ their pseudoscalar counterparts. We note that the semi-leptonic decay $H^{*}\to He^{+}e^{-}$ proceeds through both $H^{*}\to H\gamma^{*}\to He^{+}e^{-}$ and $H^{*}\to HX\to He^{+}e^{-}$. The $H^{*}\to H\gamma^{*}$ form factor is essentially $\langle H|J^{\rm em}_{\mu}|H^{*}\rangle$ with $J^{\rm em}_{\mu}=ee_{Q}\bar{Q}\gamma_{\mu}Q+ee_{q}\bar{q}\gamma_{\mu}q$. The evaluation of this matrix element was discussed in Refs. Cho:1992nt ; Amundson:1992yp ; Cheng:1992xi ; Colangelo:1993zq . It is well known that $\langle H|\bar{Q}\gamma_{\mu}Q|H^{*}\rangle$ is fixed by heavy-quark (HQ) symmetry Isgur-Wise while various approaches are employed to compute $\langle H|\bar{q}\gamma_{\mu}q|H^{*}\rangle$. Here we apply the model of vector meson dominance (VMD) BKY proposed in Ref. Colangelo:1993zq for computing the matrix element of light quark current. This choice enables the comparison of our result with that of Ref. Castro:2021gdf which also employs VMD model for computing $\langle H|\bar{q}\gamma_{\mu}q|H^{*}\rangle$. For the $H^{*}\to HX$ transition, the matrix element $\langle H|J^{X}_{\mu}|H^{*}\rangle$ with $J^{X}_{\mu}=e\varepsilon_{Q}\bar{Q}\gamma_{\mu}Q+e\varepsilon_{q}\bar{q}\gamma_{\mu}q$ can be easily inferred from $\langle H|J^{\rm em}_{\mu}|H^{*}\rangle$. We shall review the details of calculating $R_{D_{s}^{*+}}$ and correct the error made in Ref. Castro:2021gdf . In the same section, we shall also discuss the implication of $R_{D^{*0}}$ on values of $\varepsilon_{c}$ and $\varepsilon_{u}$. It will be demonstrated that the conditions $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{d}=\varepsilon_{s}$ should be relaxed in order that the data of $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ can be simultaneously accounted for. In Sec. III, we present the fitting to the couplings $\varepsilon_{u}$, $\varepsilon_{s}$, and $\varepsilon_{c}$ with the data of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, $\phi\rightarrow\eta\gamma$, and that of $R_{D_{s}^{*}}$ and $R_{D^{*0}}$. Sec. IV addresses the implications of our fitting and the outlook.
II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
We consider the $1^{-}\rightarrow 0^{-}$ heavy meson transitions $H^{*}\rightarrow He^{+}e^{-}$. The Lorentz-invariant matrix elements, accounting for the contribution by the intermediate photon and X17 boson, are denoted by $-i\mathcal{M}^{\gamma}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ and $-i\mathcal{M}^{X}\left(H^{*}\rightarrow He^{+}e^{-}\right)$, respectively. We have
$-i\mathcal{M}^{\gamma}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ $=$ $T^{\gamma}_{\mu}\frac{-ig^{\mu\nu}}{q^{2}+i\epsilon}(-ie)\bar{u}\left(p_{-}\right)\gamma_{v}v\left(p_{+}\right),$
$-i\mathcal{M}^{X}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ $=$ $T^{X}_{\mu}\frac{i\left(-g^{\mu\nu}+q^{\mu}q^{v}/m_{X}^{2}\right)}{q^{2}-m_{X}^{2}+im_{X}\Gamma_{X}}(-ie\varepsilon_{e})\bar{u}\left(p_{-}\right)\gamma_{v}v\left(p_{+}\right),$ (2)
where $e\varepsilon_{e}$ is the coupling of X17 boson to $e^{\pm}$, $\Gamma_{X}\equiv\Gamma(X\to e^{+}e^{-})=\frac{e^{2}\varepsilon_{e}^{2}m_{X}}{12\pi}\left(1+\frac{2m_{e}^{2}}{m_{X}^{2}}\right)\sqrt{1-\frac{4m_{e}^{2}}{m_{X}^{2}}}$, $T^{\gamma,X}_{\mu}$ are the $H^{*}\left(p_{H^{*}},\epsilon_{H^{*}}\right)\to H\left(p_{H}\right)\gamma^{*}(q)$ and $H^{*}\left(p_{H^{*}},\epsilon_{H^{*}}\right)\to H\left(p_{H}\right)X(q)$ decay amplitudes, respectively, and $p_{-}$ and $p_{+}$ are the momenta of final state electron and positron, respectively. By Lorentz invariance, we have
$T^{\gamma}_{\mu}$ $=$ $\langle H(p_{H})|J^{\rm em}_{\mu}|H^{*}(p_{H^{*}},\epsilon_{H^{*}})\rangle=ieF_{H^{*}H\gamma}(q^{2})\epsilon_{\mu\rho\alpha\beta}\epsilon_{H^{*}}^{\rho}p_{H^{*}}^{\alpha}p_{H}^{\beta},$
$T^{X}_{\mu}$ $=$ $\langle H(p_{H})|J^{X}_{\mu}|H^{*}(p_{H^{*}},\epsilon_{H^{*}})\rangle=ieF_{H^{*}HX}(q^{2})\epsilon_{\mu\rho\alpha\beta}\epsilon_{H^{*}}^{\rho}p_{H^{*}}^{\alpha}p_{H}^{\beta},$ (3)
with $q=p_{H^{*}}-p_{H}$. The differential decay rate for $H^{*}\rightarrow He^{+}e^{-}$ mediated by $\gamma$ is given by
$\dfrac{d\Gamma^{\gamma}(H^{\ast}\rightarrow H~{}e^{+}e^{-})}{dq^{2}}=\dfrac{\alpha_{\rm EM}^{2}}{72\pi}F_{H^{\ast}H\gamma}^{2}(q^{2})\dfrac{1}{q^{2}}\left(1+\dfrac{2m_{e}^{2}}{q^{2}}\right)\sqrt{1-\dfrac{4m_{e}^{2}}{q^{2}}}\dfrac{\lambda^{3/2}(m_{H^{\ast}}^{2},m_{H}^{2},q^{2})}{m_{H^{\ast}}^{3}},$ (4)
where $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$. Similarly, the differential decay rate mediated by X17 boson is given by
$\dfrac{d\Gamma^{X}(H^{\ast}\rightarrow H~{}e^{+}e^{-})}{dq^{2}}$ $=$ $\dfrac{\alpha^{2}_{\rm EM}\varepsilon_{e}^{2}}{72\pi}F_{H^{\ast}HX}^{2}(q^{2})\dfrac{q^{2}}{(q^{2}-m_{X}^{2})^{2}+m_{X}^{2}\Gamma_{X}^{2}}\left(1+\dfrac{2m_{e}^{2}}{q^{2}}\right)\sqrt{1-\dfrac{4m_{e}^{2}}{q^{2}}}$ (5)
$\times$ $\dfrac{\lambda^{3/2}(m_{H^{\ast}}^{2},m_{H}^{2},q^{2})}{m_{H^{\ast}}^{3}}.$
To compare with the data, we calculate the ratio $R_{H^{*}}\equiv\Gamma(H^{*}\rightarrow He^{+}e^{-})/\Gamma(H^{*}\rightarrow H\gamma)$ with the numerator containing contributions mediated by $\gamma$ and X17 boson, respectively. To separate these two contributions, we write $R_{H^{*}}=R^{\gamma}_{H^{*}}+R^{X}_{H^{*}}$. Following Eq. (4) and using $\Gamma\left(H^{*}\rightarrow H\gamma\right)=\alpha_{\rm EM}F_{H^{*}H\gamma}^{2}(0)p_{\gamma}^{3}/3$, it is easy to show that
$R_{H^{*}}^{\gamma}=\int_{q_{\rm min}^{2}}^{q_{\rm max}^{2}}\dfrac{\alpha_{\rm EM}}{3\pi}\dfrac{F_{H^{\ast}H\gamma}^{2}(q^{2})}{F_{H^{\ast}H\gamma}^{2}(0)}\dfrac{1}{q^{2}}\left(1+\dfrac{2m_{e}^{2}}{q^{2}}\right)\sqrt{1-\dfrac{4m_{e}^{2}}{q^{2}}}\dfrac{\lambda^{3/2}(m_{H^{\ast}}^{2},m_{H}^{2},q^{2})}{(m_{H^{\ast}}^{2}-m_{H}^{2})^{3}}\mbox{d}q^{2},$ (6)
where $q^{2}_{\rm min}=(2m_{e})^{2}$ and $q^{2}_{\rm max}=\left(m_{H^{*}}-m_{H}\right)^{2}$. Similarly, we apply Eq. (5) to obtain
$R_{H^{*}}^{X}$ $=$ $\int_{q_{\rm min}^{2}}^{q_{\rm max}^{2}}\dfrac{\alpha_{\rm EM}\varepsilon_{e}^{2}}{3\pi}\dfrac{F_{H^{\ast}HX}^{2}(q^{2})}{F_{H^{\ast}H\gamma}^{2}(0)}\dfrac{q^{2}}{(q^{2}-m_{X}^{2})^{2}+m_{X}^{2}\Gamma_{X}^{2}}\left(1+\dfrac{2m_{e}^{2}}{q^{2}}\right)\sqrt{1-\dfrac{4m_{e}^{2}}{q^{2}}}$ (7)
$\times$ $\dfrac{\lambda^{3/2}(m_{H^{\ast}}^{2},m_{H}^{2},q^{2})}{(m_{H^{\ast}}^{2}-m_{H}^{2})^{3}}\mbox{d}q^{2}.$
The above equation can be simplified by applying the narrow width approximation,
$\frac{1}{\left(q^{2}-m_{X}^{2}\right)^{2}+m_{X}^{2}\Gamma_{X}^{2}}=\frac{\pi}{m_{X}\Gamma_{X}}\delta\left(q^{2}-m_{X}^{2}\right).$ (8)
For the case of $D^{*}$ meson decays, the form factors $F_{H^{*}H\gamma}(q)$ and $F_{H^{*}HX}(q)$ were calculated in Castro:2021gdf .
Figure 1: The experimentally measured as well as theoretically predicted values for $R_{D_{s}^{*+}}$. Various theoretical predictions are distinguished by different values taken for $\varepsilon_{c}$ and $\varepsilon_{s}$. VMD__\__ 2021 stands for the calculation of Castro:2021gdf using VMD approach and employing $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=7.4\times 10^{-3}$ extracted by Feng:2016jff ; Fornal:2017msy . VMD__\__ 2021 (corrected result) refers to our recalculation of $R_{D_{s}^{*+}}$ using the same values of $\varepsilon_{c}$ and $\varepsilon_{s}$. VMD__\__ 2023 denoted in blue represents our result of using $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays, and that denoted in black is obtained by employing $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The SM prediction is denoted as SM__\__ Photon.
We note that the values of $\varepsilon_{Q}$ and $\varepsilon_{q}$ significantly affect $R_{H^{*}}^{X}$. Fig. 1 presents the experimentally measured value of $R_{D^{*+}_{s}}$ as well as the corresponding theoretical predictions with and without contributions from the X17 boson. The measured value of $R_{D^{*+}_{s}}$ is marked in red while the SM prediction of this ratio is marked in yellow, which considers only the photon contribution and gives $R_{D^{*+}_{s}}^{\gamma}=6.7\times 10^{-3}$. This is consistent with the experimental data, $R_{D^{*+}_{s}}^{\text{data}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . With the contribution of X17 boson included, VMD__\__ 2023 contains two cases. The first one, denoted in blue, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays. The second case, denoted in black, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The uncertainties for theoretical predictions arise from the uncertainty of coupling parameter $\lambda$ in the interaction vertex $\lambda H^{*}HV$ with $V=\rho^{0},\ \omega,\ \phi$. This parameter appears in the form factors $F_{H^{*}HX}(q^{2})$ and $F_{H^{*}H\gamma}(q^{2})$ and takes the value Colangelo:1993zq
$\lambda=(-0.289\pm 0.016)\ {\rm GeV}^{-1}.$ (9)
Theoretical predictions denoted by blue and black line segments in Fig. 1 are consistent with the experimental data. On the other hand, $R_{D_{s}^{*}}$ calculated with $\varepsilon_{u}=$$\pm 3.7\times 10^{-3},\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff yields $(3.3\pm 1.3)\times 10^{-2}$, which significantly deviates from the data. Our result disagrees with the calculation by Castro:2021gdf which gives $R_{D_{s}^{*+}}=(9.82\pm 0.60)\times 10^{-3}$. The latter turns out to be not far from the measurement $R_{D^{*+}_{s}}^{\text{data}}$.
Table 1: The values $R_{H^{*}}^{\gamma}$ and $R_{H^{*}}^{X}$ calculated with VMD model and the coupling values $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in ${}^{8}{\rm Be}^{*}$ decays considered Denton:2023gat .
Decay mode $R_{H^{*}}^{\gamma}$ $R_{H^{*}}^{X}$ $R_{H^{*}}$ Experiment
$D^{\ast 0}\rightarrow D^{0}e^{+}e^{-}$ $6.7\times 10^{-3}$ $5.5\times 10^{-7}$ $6.7\times 10^{-3}$ $(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq
$D^{\ast+}\rightarrow D^{+}e^{+}e^{-}$ $6.6\times 10^{-3}$ $(1.8\pm 0.1)\times 10^{-4}$ $6.8\times 10^{-3}$
$D_{s}^{\ast+}\rightarrow D_{s}^{+}e^{+}e^{-}$ $6.7\times 10^{-3}$ $(5.7\pm 3.0)\times 10^{-3}$ $(1.2\pm 0.3)\times 10^{-2}$ $(7.2_{-1.6}^{+1.8})\times 10^{-3}$ CLEO:2011mla
We consider a similar analysis for the decay $D^{*0}\rightarrow D^{0}e^{+}e^{-}$. The experimental measurement gives $R^{\text{exp}}_{D^{*0}}=(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq and it is marked in red in Fig. 2. The SM_photon scenario yields $R_{D^{*0}}^{\gamma}=6.7\times 10^{-3}$. The full contributions $R_{D^{*0}}\equiv R_{D^{*0}}^{\gamma}+R_{D^{*0}}^{X}$ corresponding to different ranges of new physics couplings are labeled as VMD_2021 and VMD_2023, respectively. Table 1 presents results of $R_{H^{*}}$ arising from intermediate photon and X17 boson contributions for three $D^{*}$ meson decay modes within the VMD model, employing $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in ${}^{8}{\rm Be}^{*}$ decays considered Denton:2023gat . Considering isospin mixing and breaking effects in ${}^{8}{\rm Be}^{*}$ decays, i.e., $\varepsilon_{u}=\varepsilon_{c}=\pm 9.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.5\times 10^{-3}$, the corresponding $R_{H^{*}}$ values are $6.7\times 10^{-3}$, $6.8\times 10^{-3}$, and $(1.0\pm 0.4)\times 10^{-2}$ for $D^{*0}$, $D^{*+}$, and $D^{*+}_{s}$ decays, respectively.
From Fig. 2, one can see that the observed $R_{D^{*0}}$ is much larger than every theoretical prediction. This implies that the intermediate X17 boson does not significantly enhance $D^{*0}\to D^{0}e^{+}e^{-}$ decay to match with the experimental data. The most straightforward remedy for the situation is by increasing the value of $\varepsilon_{c}$. However, an overly large $\varepsilon_{c}$ can easily disrupt the agreement between theory and experiment on $R_{D^{*}_{s}}$. To circumvent this problem, the value of $\varepsilon_{s}$ must also be adjusted accordingly. It is clear that we have to remove the assumption of $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$.
Figure 2: The experimentally measured and theoretically predicted values for $R_{D_{0}^{*}}$. VMD__\__ 2021 is the result obtained by VMD approach where $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff ; Fornal:2017msy . The result of VMD__\__ 2023 denoted in blue is calculated by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$, and that of VMD__\__ 2023 denoted in black is obtained by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. The SM prediction is denoted as SM__\__ Photon.
We observe that the above adjustments of $\varepsilon_{c}$ and $\varepsilon_{s}$ naturally affect the decays of Charmonium ($c\bar{c}$) and $\phi$ ($s\bar{s}$) mesons. Hence the determination of these coupling parameters should involve the data of such decay modes.
III X17 couplings to charm and strange quarks
This section aims to incorporate the decay channels $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$and $\phi\rightarrow\eta e^{+}e^{-}$ for determining X17 couplings to charm and strange quarks. We use VMD approach to calculate the normalized transition form factor (TFF) $\mathcal{F}_{H^{*}HV}\left(q^{2}\right)$
$\mathcal{F}_{H^{*}HV}\left(q^{2}\right)=\frac{F_{H^{*}HV}\left(q^{2}\right)}{F_{H^{*}H\gamma}(0)},$ (10)
with $V=\gamma,X$. From Eq. (3), we have
$ee_{c}\langle\eta_{c}(p_{f})|\bar{c}\gamma_{\mu}c|\psi(2S)(p_{i},\epsilon_{i})\rangle$ $=$ $ieF_{\psi(2S)\eta_{c}\gamma}(q^{2})\epsilon_{\mu\rho\alpha\beta}\epsilon_{i}^{\rho}p_{i}^{\alpha}p_{f}^{\beta},$
$e\varepsilon_{c}\langle\eta_{c}(p_{f})|\bar{c}\gamma_{\mu}c|\psi(2S)(p_{i},\epsilon_{i})\rangle$ $=$ $ieF_{\psi(2S)\eta_{c}X}(q^{2})\epsilon_{\mu\rho\alpha\beta}\epsilon_{i}^{\rho}p_{i}^{\alpha}p_{f}^{\beta},$ (11)
with $q=p_{i}-p_{f}$. It is clear that
$\mathcal{F}_{\psi(2S)\eta_{c}X}\left(q^{2}\right)=\left(\frac{\varepsilon_{c}}{e_{c}}\right)\times\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}\right)=3\varepsilon_{c}\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}\right)/2.$ (12)
$\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}\right)$ is the normalized TFF for a transition between a $\psi(3686)$ and $\eta_{c}$ meson mediated by a virtual photon with the momentum $q$. It has been characterized by the pole approximation as Gu:2019qwo
$\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}\right)=\frac{1}{1-q^{2}/\Lambda_{\psi(2S)\eta_{c}}^{2}}.$ (13)
The VMD assumption has been phenomenologically successful in explaining TFF behaviors in similar decays. The effective pole mass refers to the mass of the vector meson resonance that is close to the energy scale of the decaying particle. In the analysis of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\Lambda_{\psi(2S)\eta_{c}}$ corresponds to the mass of a higher excited state, $m_{\psi(3770)}=(3773.7\pm 0.4)\,\mathrm{MeV/c^{2}}$ Gu:2019qwo . This framework has provided a coherent explanation for the experimental observations in the Charmonium decays. From Eq. (7), the ratio $R_{\psi(2S)}^{X}$ between the decay rate of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$ decay rate and that of $\psi(2S)\rightarrow\eta_{c}\gamma$ is given by
$R_{\psi(2S)}^{X}=9\varepsilon_{c}^{2}/4.$ (14)
This is an approximated result in the limit $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. The measurement by BESIII gives ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}\right)=$ $\left(3.77\pm 0.40_{\text{stat }}\pm 0.18_{\text{syst }}\right)\times 10^{-5}$ BESIII:2022jde . Utilizing the branching ratio ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}\gamma\right)=(3.6\pm 0.5)\times 10^{-3}$ ParticleDataGroup:2024cfk , one obtains the ratio $R^{\rm data}_{\psi(2S)}=(1.1\pm 0.2)\times 10^{-2}$ as reported in BESIII:2022jde as well. The theoretical prediction for the photon-mediated decay rate given by the VMD model is $R_{\psi(2S)}^{\gamma}=8.9\times 10^{-3}$. The tiny difference between $R_{\psi(2S)}^{\gamma}$ and $R^{\rm data}_{\psi(2S)}$ gives a strong constraint to $R_{\psi(2S)}^{X}$, i.e., the coupling parameter $\varepsilon_{c}$.
The decay $\phi\rightarrow\eta e^{+}e^{-}$is a transition from a pure $\bar{s}s$ state to a mixed state involving up ($\bar{u}u$), down ($\bar{d}d$), and strange ($\bar{s}s$) quark pairs, emitting an electron-positron pair in the process. The measurements by SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk yield the branching ratio for $\phi\rightarrow\eta e^{+}e^{-}$ as $\left(1.19\pm 0.19_{\text{stat. }}\pm\right.$ $\left.0.07_{\text{syst. }}\right)\times 10^{-4},\left(1.14\pm 0.10_{\text{stat. }}\pm 0.06_{\text{syst. }}\right)\times 10^{-4}$ and $\left(1.075\pm 0.007_{\text{stat. }}\pm 0.038_{\text{syst. }}\right)\times 10^{-4}$, respectively. The VMD approach is utilized to estimate the lepton mass spectrum for the electromagnetic (EM) Dalitz decay $\phi\rightarrow\eta e^{+}e^{-}$. Following the similar approach as that of Eq. (11), we obtain the relation
$\mathcal{F}_{\phi\eta X}\left(q^{2}\right)=\left(\frac{\varepsilon_{s}}{e_{s}}\right)\times\mathcal{F}_{\phi\eta\gamma}\left(q^{2}\right)=-3\varepsilon_{s}\mathcal{F}_{\phi\eta\gamma}\left(q^{2}\right).$ (15)
Once more $\mathcal{F}_{\phi\eta\gamma}$ can be expressed as a simple pole Faessler:1999de :
$\mathcal{F}_{\phi\eta\gamma}\left(q^{2}\right)=\frac{1}{1-q^{2}/\Lambda_{\phi\eta}^{2}}.$ (16)
[…]
$R_{\phi}^{X}=9\varepsilon_{s}^{2}.$ (17)
Similar to the case of $\psi(2S)$ decay, we obtain this result with the approximations $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\phi\eta\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. Given the measurements on $R_{\phi}$ from three separate experiments and the theoretical prediction, $R_{\phi}^{\gamma}=8.30\times 10^{-3}$, we are able to determine the allowed range for $R_{\phi}^{X}$.
Refer to caption Refer to caption
(a) (b)
Figure 3: The allowed ranges for (a) $\left(\varepsilon_{c},\varepsilon_{s}\right)$ and (b) $\left(\varepsilon_{c},\varepsilon_{u}\right)$ extracted by the data of $D^{*+}_{s}$, $D^{*0}$, $\psi(2S)$ and $\phi$ decays.
The $1\sigma$ parameter range allowed by the ratio between $\Gamma(\psi(2S)\to\eta_{c}e^{+}e^{-})$ BESIII:2022jde and $\Gamma(\psi(2S)\to\eta_{c}\gamma)$ is denoted by the pink area in Fig. 3. The left panel depicts the allowed range for the $\left(\varepsilon_{c},\varepsilon_{s}\right)$ pair while the right panel depicts that for $\left(\varepsilon_{c},\varepsilon_{u}\right)$. Furthermore, the left panel of Fig. 3 also shows $1\sigma$ ranges for $\varepsilon_{s}$ from the data of SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk on $\phi\to\eta e^{+}e^{-}$, denoted by yellow, gray, and purple areas, respectively. The blue band is the allowed range for $\left(\varepsilon_{c},\varepsilon_{s}\right)$ by the data of $D^{*+}_{s}\to D^{+}_{s}e^{+}e^{-}$ CLEO:2011mla . On the right panel of Fig. 3, the $1\sigma$ allowed range for $\left(\varepsilon_{c},\varepsilon_{u}\right)$ by the data of $D^{*0}\to D^{0}e^{+}e^{-}$ is denoted by the green areas. Finally the extracted range for $|\varepsilon_{u}|$ from Refs. Feng:2016jff and  Denton:2023gat are denoted by black dashed curves and red band, respectively.
Table 2: The measured values and the corresponding Standard Model (SM) predictions of four decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons, respectively.
Decay Ratio Measured Value SM Prediction
$\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ $(11\pm 1)\times 10^{-3}$ $6.7\times 10^{-3}$
$\Gamma(D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-})/\Gamma(D_{s}^{*+}\rightarrow D_{s}^{+}\gamma)$ $\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ $6.7\times 10^{-3}$
$\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ $(11\pm 2)\times 10^{-3}$ $8.9\times 10^{-3}$
$\Gamma(\phi\rightarrow\eta e^{+}e^{-})/\Gamma(\phi\rightarrow\eta\gamma)$ $(8.31\pm 0.33)\times 10^{-3}$ $8.30\times 10^{-3}$
We employ the $\chi^{2}$ statistical test for fitting the coupling parameters, $\varepsilon_{c},\varepsilon_{s}$, and $\varepsilon_{u}$ with the four measurements given in Table 2. For $\phi\to\eta e^{+}e^{-}$ decay, we only employ the data by KLOE-2 KLOE-2:2014hfk since its uncertainty is much smaller than those of earlier measurements. We minimize the $\chi^{2}$ defined as
$\chi^{2}=\sum_{i=1}^{4}\frac{\left(R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\varepsilon_{u}\right)-R_{i}^{\rm ob}\right)^{2}}{\sigma_{i}^{2}}+\frac{\left(\lambda-\lambda_{\rm best\,fit}\right)^{2}}{\sigma_{\lambda}^{2}},$ (18)
where $R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\varepsilon_{u}\right)$ denotes the theoretically predicted value for the $i$-th decay ratio, which depends on the parameters $\varepsilon_{c}$, $\varepsilon_{s}$, and $\varepsilon_{u}$. $R_{i}^{\rm ob}$ refers to the observed value for the $i$-th decay ratio with $\sigma_{i}$ the corresponding uncertainty for the measurement. The parameter $\lambda$ determines the coupling strength of the heavy meson to light vector mesons, i.e., it appears in the coupling vertex $\lambda H^{*}HV$. From Eq. (9), we have $\lambda_{\rm best\,fit}=-0.289$ GeV-1 and $\sigma_{\lambda}=0.016$ GeV-1. Our fitting uses the measured decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons as outlined in Table 2.
[…]
The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=3\times 10^{-4}$ and $\left|\varepsilon_{s}\right|=4\times 10^{-5}$ while there are two best-fit solutions for $\left|\varepsilon_{u}\right|$, which are $4.7\times 10^{-2}$ and $5.0\times 10^{-2}$, respectively. The fitting results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. Up to the $3\sigma$ range, the value of $\left|\varepsilon_{u}\right|$ is of the order $10^{-2}$, which is notably larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ Denton:2023gat . In other words, the value of $\left|\varepsilon_{u}\right|$ extracted from the decays of $D^{*}$, $\psi(2S)$, and $\phi$ mesons is in a serious tension with that extracted by ATOMKI measurements. The significant increase of $\left|\varepsilon_{u}\right|$ in this case can be attributed to the anomalously large branching ratio of $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ measured by by BESIII BESIII:2021vyq as illustrated by Fig. 2 and the right panel of Fig. 3. We remark that only the decay of $D^{*0}$ among those of $D_{s}^{*+}$, $D^{*0}$, $\psi(2S)$, and $\phi$ mesons involve X17 couplings to the first generation of quarks, i.e., $\varepsilon_{u}$ specifically.
IV Conclusions
The confirmation of couplings between X17 boson and quarks remains a topic of interest to date. The hypothesis of X17 boson has continuously received attentions by the data of anomalous 8Be, 4He, and 12C decays over time. The accumulation of data has shifted the most favored coupling strengths from $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ Feng:2016jff to $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ without isospin effects, and $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ with both isospin mixing and breaking effects considered Denton:2023gat . The persistent decay anomalies outlined above have motivated the testing of X17 hypothesis in the decays of heavy-flavor mesons such as $D^{*}$ and $B^{*}$ mesons with the assumption $\varepsilon_{b}=\varepsilon_{s}=\varepsilon_{d}$ and $\varepsilon_{c}=\varepsilon_{u}$ Castro:2021gdf . We have revisited this study by focusing on the comparisons of theoretical predictions (with X17 boson mediated effects included) to decay ratios $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ with experimental data CLEO:2011mla ; BESIII:2021vyq . For the case of $R_{D_{s}^{*+}}$, we have found that only the updated values of $\varepsilon_{u,d}$ are consistent with the measurement while the earlier extracted parameter values $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ predict a $R_{D_{s}^{*+}}$ much greater than the measurement (see Fig. 1). Our finding corrects an error by Ref. Castro:2021gdf which concluded that the prediction by $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ is consistent with the measurement.
To accommodate the measured $R_{D^{*0}}$, one requires a much larger $\left|\varepsilon_{u}\right|$ under the assumption $\varepsilon_{c}=\varepsilon_{u}$. On the other hand, raising the magnitude of $\varepsilon_{c}$ would affect $R_{D_{s}^{*+}}$ unless the magnitude of $\varepsilon_{s}$ is increased as well since $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ as indicated by Fig. 3 (left panel). Since $\varepsilon_{s}$ already departs from $\varepsilon_{d}$ in this case, we then also treat $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent parameters in our analysis. To determine the favored ranges for $\varepsilon_{u}$, $\varepsilon_{c}$ and $\varepsilon_{s}$, we have performed a fitting to the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}$, and $R_{\phi}$. The results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=3\times 10^{-4}$, $\left|\varepsilon_{s}\right|=4\times 10^{-5}$, and two solutions for $\left|\varepsilon_{u}\right|$, which are $4.7\times 10^{-2}$ and $5.0\times 10^{-2}$, respectively.
It is noteworthy that the value of $\left|\varepsilon_{u}\right|$ remains of the order $10^{-2}$ up to the $3\sigma$ range. Remarkably, this value is much larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|=(0.5-0.9)\times 10^{-3}$ Denton:2023gat . Such a tension is caused by the large deviation of the measured $R_{D^{*0}}$ from the Standard Model prediction as shown in Table 2. This is also reflected by the right panel of Fig. 3, where both green bands significantly deviate from the line $\varepsilon_{u}=0$. In conclusion, we have studied effects of the proposed X17 boson to the decays of $D^{*}$, $\psi(2S)$, and $\phi$ mesons, which involve X17 boson couplings to both the first and second generations of quarks. The above-mentioned tension on the value of $\varepsilon_{u}$ deserves further scrutiny. We note that the decay $D^{*\pm}\to D^{\pm}e^{+}e^{-}$ remains to be measured. Furthermore the huge deviation of the measured $D^{*0}\to D^{0}e^{+}e^{-}$ branching ratio to its Standard Model predicted value awaits confirmations from updated measurements.
Acknowledgements
This work is supported by National Science and Technology Council, Taiwan, under Grant No. NSTC 112-2112-M-A49-017.